OpenStudy (anonymous):

solve the system of equations using elimination type an ordered pair x+y=9 -x+2y=9

OpenStudy (anonymous):

-x=y-9 so you substitute and get y-9+2y=9 3y=18 y=6 from the top -x=6-9 -x=-3 x=3

OpenStudy (anonymous):

what is the ordered pair